Upstream, downstream. Junior’s boat will go 15 miles per hour in still water. If he can go 12 miles downstream in the same amount of time as it takes to go 9 miles upstream, then what is the speed of the current? D R T Downstream 12 x 12/x Upstream 9 x 9/x The speed of current is 1/5 mph. On the right track or totally off? So far..... 12/x - 9/x = 15 ; 3/x = 15 ; 3 = 15x ; 3/15 = x ; 1/5 = x ; Am I on the right track?...Thank you

Answer by khwang(438)

Let the speed of the current be x mph.
[Note x should be less than 15 mph, for otherwise theboat cannot move
when gotupstream.]
Downstream speed = x + 15. (sum of the speeds)
Upstream speed = 15-x, (difference of the speeds)
Now we have 12/(x+15) = 9/(15-x) or
4(15-x) = 3(x+15)
so 7x = 15 , x = 15/7 mph.
You did completely wrong , try to think about it.
Kenny